OpenStudy (anonymous):

can anyone help me step by step please!! 3(-9n + 5) + -4 = -16

OpenStudy (anonymous):

3(-9n + 5) + -4 = -16 distribut the 3 -27n+15 +-4=-16 combine like terms -27n+11=-16 subtract 11 from both sides -27n=-27 divide both sides from -27 n=1

OpenStudy (anonymous):

okay? good? okay

OpenStudy (anonymous):

@Tahmina thanks

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

u welcome :)

OpenStudy (anonymous):

anymore help? just post it up here okay?